### v3

# PET-SQL: A Prompt-enhanced Two-stage Text-to-SQL Framework with Cross-consistency

Abstract
Recent advancements in Text-to-SQL (Text2SQL) emphasize stimulating the large language models (LLM) on in-context learning, achieving significant results. Nevertheless, they face challenges when dealing with verbose database information and complex user intentions. This paper presents a two-stage framework to enhance the performance of current LLM-based natural language to SQL systems. We first introduce a novel prompt representation, called reference-enhanced representation, which includes schema information and randomly sampled cell values from tables to instruct LLMs in generating SQL queries. Then, in the first stage, question-SQL pairs are retrieved as few-shot demonstrations, prompting the LLM to generate a preliminary SQL (PreSQL). After that, the mentioned entities in PreSQL are parsed to conduct schema linking, which can significantly compact the useful information. In the second stage, with the linked schema, we simplify the prompt’s schema information and instruct the LLM to produce the final SQL. Finally, as the post-refinement module, we propose using cross-consistency across different LLMs rather than self-consistency within a particular LLM. Our methods achieve new SOTA results on the Spider benchmark, with an execution accuracy of 87.6%. The codes are released on https://github.com/zhshLii/PETSQL
[…]
1 Introduction
Text-to-SQL (Text2SQL) assists individuals in converting natural language questions (i.e., text) into structure query language (SQL) queries. The task is helpful for question-answering and information-retrieval applications such as traffic and business intelligence [1; 2], as it lowers the requirements for expertise and allows non-professional participants to interact with databases in natural language.
Recent advancements in Text2SQL methods emphasize stimulating the performance of large language models (LLM) on in-context learning with elaborate prompt strategies [3; 4; 5]. Couple with it, typical post-refinements that rely exclusively on executing results or exceptions are proposed, such as self-debugging [1; 6] and self-consistency [4]. However, these approaches encounter difficulties when handling extensive databases across domains and intricate user intentions. First, during the design of the prompt, recent studies focus on the representations/processing of schema information, while lacking a priori attention to cell values format in tables. This relates to the handling of conditionals. For example, the use of conditions ‘sex=‘Male’’, ‘sex=‘male’’, or ‘sex=‘M’’ will produce different results when querying for the gender. In addition, special instructions in the prompt need to be further optimized. Second, to simplify the verbose schema and thus reduce the heavy context, the canonical schema-linking module is usually an independent plugin that instructs LLM to yield question-aware tables and columns for simplification. It consumes the LLM budget once but only produces limited information (i.e., related table and column names). Moreover, the outcomes may lack conciseness, and for coding LLMs like CodeLlama [7], generating non-code (plain) text based on instructions is not their forte. Third, the gains of current post-refinements are not significant. For example, self-debugging is inaccessible if the SQLs generated by a strong LLM have semantic ambiguity rather than syntax errors (our experience with multiple LLMs, especially GPT4). In addition, the investment and production in self-consistency of one LLM are disproportionate, as it calls the same LLM several times (time-consuming and yields many times the cost) at a higher temperature, which does not fundamentally diversify the SQL outputs. The higher temperature may generate more hallucinations, potentially harming the performance of the LLM [8].
In light of the above challenges, in this paper, we present the Prompt-Enhanced Two-stage text-to-SQL framework with cross-consistency (PET-SQL for short), that stimulates the performance of current LLMs without the need to train specific models. Concretely, we first elaborate on a general and effective representation of the prompt, termed reference-enhanced representation. Besides the schema information, the cell values in tables are randomly sampled and added to provide oversight of the value format. We also instruct the LLM to minimize SQL execution time while ensuring correctness. Then, based on the prompt, in the first stage, following the question de-semanticization setting in [9], several question-SQL pairs are retrieved as few-shot demonstrations and prompt LLM to generate the preliminary SQL (PreSQL). After that, the mentioned tables in PreSQL are parsed, which serves as the rationale for schema linking. In the second stage, with the linked schema, we simplify the schema information in the prompt and ask the LLM to yield the final SQL (FinSQL). Finally, we propose to use the cross-consistency of different LLMs rather than the self-consistency of a particular LLM, since we believe that diversity across different LLMs is more beneficial than the repeated calling for an LLM under high temperatures. Overall, the method achieves state-of-the-art in the Spider [10] leaderboard with 87.6% execution accuracy, which achieves a 1% improvement against the DAIL-SQL [4].
The main contributions of the work are summarized as follows:
  • A one-for-all elaborated prompt is designed and its effectiveness is systemically evaluated under various LLMs. Besides the schema information, the cell values in tables and magic instructions are attended to as the prior knowledge in the prompt.
  • We explore the PreSQL-based schema linking instead of prompting the LLM to directly generate table/column names. We believe it can yield more concise results and is suitable for coding LLM. Coupled with linked schema, we present to organize the simplified prompt and then feed it into the LLM again.
  • The cross-consistency is proposed to utilize the diversity across different LLMs. The PreSQL is also re-used to vote in this module. Overall, by voting across various LLMs, the final predicted SQL achieves 87.6% execution accuracy in the Spider benchmark.
2 Related Work
Learning-based agents. With the advancement of language models in previous years, multiple methods have been proposed for the Text2SQL task. These early approaches focus on pattern matching between natural language (NL) and SQL statements [11; 12; 13]. Some studies employ relation-aware self-attention mechanisms as encoders to learn representations of questions and schemas. These methods then utilize grammar-based decoders to generate SQL as abstract syntax trees or employ sketch-based decoders to obtain SQL by filling in slots [14; 15]. Subsequently, the advent of pre-trained language models and the fine-tuning paradigm significantly influences these methods [16]. Numerous studies utilize standard sequence-to-sequence models with transformer architectures [17] to translate NL questions into SQL queries in an end-to-end manner [18; 19].
LLM-based in-context learning methods. The development of LLMs has catalyzed substantial transformations in this field. These methods leverage the in-context learning, semantic understanding, and reasoning capabilities of LLMs [20; 21; 22], continuously pushing the boundaries of performance on various evaluation benchmarks, such as Spider [10] and BIRD [23]. For example, C3 [3] achieves a breakthrough by leveraging ChatGPT’s zero-shot learning capability. DIN-SQL  [5] decomposes the task into multiple components, including schema linking, difficulty classification, and SQL generation. This methodology effectively reduces the complexity of decision-making, enabling LLMs to generate more precise SQL statements. DAIL-SQL [4] further enhances the capabilities of LLMs through in-context learning and supervised fine-tuning schemes. In addition, several other studies incorporate advanced reasoning methods, such as chain-of-thought [24] and self-reflexion [25], to enhance the capabilities of LLMs and improve their performance on Text2SQL tasks [26; 27].
3 Methodology
According to a natural language question $Q$ on a database $D$, the objective of LLM-based Text2SQL is to translate $Q$ into an executable SQL query s{s}. The likelihood of an LLM $\mathcal{M}$ generating a SQL query s{s} can be formally defined as the conditional probability distribution
$\mathbb{P}_{\mathcal{M}}(s|\mathcal{P}({Q},\mathcal{I}(D)))=\prod_{i=1}^{|s|}\mathbb{P}_{\mathcal{M}}(s_{i}|\mathcal{P}(Q,\mathcal{I}(D)),s_{<i}),$ (1)
where $\mathcal{I}(D)$ means the description or information of $D$, such as data definition language. $\mathcal{P}$ is the prompt templete, and $|s|$ is the length of $s$. $s_{i}$ and $s_{<i}$ denotes the $i$-th token and the prefix of $s_{i}$, respectively.
In what follows, we will detail the PET-SQL framework. As shown in Fig. 1, its specifications can be summarized by four points: (1) an elaborated prompt that harnesses the customized instructions, basic database information, and samples in the stored tables (2) instructing the LLM to generate PreSQL, in which some demonstrations are selected from pools using a question similarity-based strategy and then prefixed to the prompt as the few-shot in-context (3) finding question-related tables (schema linking) based on the PreSQL and prompt LLM to yield FinSQL by the linked schema (4) ensuring consistency in the predicted results across multiple LLMs.
Figure 1: The overview of the proposed PET-SQL framework
3.1 Prompting
When instructing LLMs to generate SQL queries, the usage of styles or templates for prompts significantly impacts LLMs’ performance. As [4] recommended, Code Representation ($CR_{p}$) and OpenAI Demonstration ($OD_{p}$) forms mixed with additional information are better choices (shown in Listing 1). To further exploit the potential of LLM, we enrich the prompt based on $OD_{p}$ which is termed reference-enhanced representation ($RE_{p}$).
1OpenAI Demonstration Format:
2### How many singers do we have?
3SELECT count(*) FROM singers
4
5Code Representation Format:
6/* How many singers do we have? */
7SELECT count(*) FROM singers
Listing 1: The simple showcases of OpenAI Demonstration and Code Representation formats.
Building upon $OD_{p}$ which consists of the task instruction, database schema, and question components, $RE_{p}$ performs three modifications: optimization rule (OR), cell value references (CV), and foreign key declarations (FK). Concretely, in the task instruction, we raise a multi-task constraint rule for LLM, that is, highlighting the LLM to “minimize SQL execution time while ensuring correctness” (line 1 in Listing 2). The rationale is not only to focus on execution accuracy but also to make LLM aware of the efficiency of SQL statements. During the generation of efficient SQL statements by LLM, redundant characters and operators can be avoided (which often result in exceptions). Ablation experiments also confirm the effectiveness of the rule. Second, three rows in each table are randomly sampled and inserted in the prompt (lines 7-11 in Listing 2). The sampled cell values as references can help LLM understand the formats and specifications of the database. It relieves the dilemma that being unsure which element to query as the condition, caused by the lack of standardization in filling out different tables. For example, all of the “M”, “Male”, and “male” are feasible representations of gender. Without prior knowledge of cell values, organizing condition statements to query the number of men in a table can be confusing. Thirdly, foreign key relations in the schema are added as suffixes in the prompt (lines 12-15 in Listing 2). It facilitates LLM to recognize the connections between the tables involved in the database, and thus improves the understanding of the user’s intent and automatically selects the appropriate connection in the queries.
It should be noted that the above-detailed prompt is in the zero-shot setting and can also be enhanced by the few-shot approach. Next, we explain how to select demonstrations as shots and equip them with the prompt and thus stimulate the LLM to yield the PreSQL for a question.
1### Answer the question by SQLite SQL query only and with no explanation. You must minimize SQL execution time while ensuring correctness.
2### Sqlite SQL tables, with their properties:
3#
4# singer(Singer_ID,Name,Country,Song_Name,Song_release_year,Age,Is_male);
5# singer_in_concert(concert_ID,Singer_ID);
6#
7### Here is some data information about database references.
8#
9# singer(Singer_ID[1,2,3],Name[Joe,Timbaland,Justin Brown],Country[Netherlands,United States,France],Song_Name[You,Dangerous,Hey Oh],Song_release_year[1992,2008,2013],Age[52,32,29],Is_male[F,T,T]);
10# singer_in_concert(concert_ID[1,1,1],Singer_ID[2,3,5]);
11#
12### Foreign key information of SQLite tables, used for table joins:
13#
14# singer_in_concert(Singer_ID) REFERENCES singer(Singer_ID);
15#
16### Question: How many singers do we have?
17### SQL:
Listing 2: An example of the proposed zero-shot reference-enhanced representation.
3.2 Question skeleton-based PreSQL generation
The goal of this module is to retrieve domain-agnostic samples from known datasets (i.e., the train set) that exhibit similar question intent as the target question. The retrieved question-SQL samples can serve as demonstrations for the in-context learning of LLM.
Following the question de-semanticization module in [9], we mask the domain-related tokens (i.e., table names, column names, and values) in questions with a special token <mask> according to the database schema, and thus obtain the question skeletons which represent the question’s intentions. Then, all the question skeletons and question-SQL pairs in the training set are constructed as a demonstration pool. Based on the semantic embeddings of question skeletons, we retrieve the top-K similar samples from the pool with the target question. Like [4], the embedding model for question skeletons is a pre-trained sentence Transformer. After that, we organize the selected demonstrations with our prompt $RE_{p}$ following [4], that is, prefixing question-SQL pairs as the few-shot contexts in $RE_{p}$ (as shown in lines 1-10 in Listing 3). Finally, the few-shot $RE_{p}$ is used to prompt the LLM and generate PreSQL.
1### Some example pairs of questions and corresponding SQL queries are provided based on similar questions:
2
3### How many farms are there?
4SELECT count(*) FROM farm
5
6### How many books are there?
7SELECT count(*) FROM book
8
9### How many actors are there?
10SELECT count(*) FROM actor
11
12\${zero-shot reference-enhanced representation}
Listing 3: Three question-SQL pairs are retrieved and prefix as a 3-shot prompt.
3.3 Schema linking and FinSQL generation
To eschew the verbose information in the schema that may obstruct the LLMs’ performance, we further conduct the schema linking (SL) which involves identifying references (table/column) to the database schema and condition values related to the natural language question. After that, the linked schema is extracted and utilized to reorganize the prompt, which is then fed into the LLM again.
The implementation of schema linking is based on a simple principle: The schema information mentioned in the PreSQL is concise and relevant to the question, and thus the table/column entities in the PreSQL statement can be directly parsed as the linking results. Rather than designing strategies such as sorting [3] or extraction [5] to make LLM output relevant database references, it is better to generate complete SQL statements directly. There are three advantages: (1) Most LLMs are pre-trained with Text2SQL or SQL-related corpus, so presumably their ability on SQL generation is stronger than the schema output (e.g., a list with table.column format) following instructions. (2) Generating SQL is more generic and applicable for coding-specific LLMs such as CodeLlama [7]. (3) The execution accuracy is theoretically monotonic. For a correctly executed PreSQL, the linked results are always correct. Therefore, the resulting FinSQL should also be correct. Wrong PreSQL may be attributed to the LLM being confused by the verbose prompt, and there is room to be correct with a concise and simplified prompt. It should be noted that column linking has poor fault tolerance and may be inapplicable to weak LLMs, so we only perform table linking at this stage.
Then, we simplify the prompt with the parsed schemas. The simplifications are shown in Listing 4. All the linked-table-irrelevant contexts in the prompt are removed, including the schema properties, database references, and foreign key declarations. Following that, the simplified prompt will be fed into LLMs to generate the FinSQL.
1### Answer the question by SQLite SQL query only and with no explanation. You must minimize SQL execution time while ensuring correctness.
2### Sqlite SQL tables, with their properties:
3#
4# singer(Singer_ID,Name,Country,Song_Name,Song_release_year,Age,Is_male);
5# singer_in_concert(concert_ID,Singer_ID);
6#
7### Here is some data information about database references.
8#
9# singer(Singer_ID[1,2,3],Name[Joe,Timbaland,Justin Brown],Country[Netherlands,United States,France],Song_Name[You,Dangerous,Hey Oh],Song_release_year[1992,2008,2013],Age[52,32,29],Is_male[F,T,T]);
10# singer_in_concert(concert_ID[1,1,1],Singer_ID[2,3,5]);
11#
12### Foreign key information of SQLite tables, used for table joins:
13#
14# singer_in_concert(Singer_ID) REFERENCES singer(Singer_ID);
15#
16### Question: How many singers do we have?
17### SQL:
Listing 4: A simplified prompt when the linked table from PreSQL is “singer”. The strikethrough (in lines 5, 10-14) indicates removing contexts in $RE_{p}$ via schema linking.
3.4 Cross consistency
In this section, we elaborate on the designed consistency module based on the executed results of SQL queries on the database engine, including two feasible implementation strategies: naive voting across several LLMs and fine-grained voting according to PreSQL complexity.
As a canonical trick, the self-consistency approach is widely employed for post-refining the executed results, which is conducted by increasing the randomness of an LLM output at high temperatures to generate diverse results and followed by majority voting among multiple executed results. However, it has been reported that high temperatures may lower the performance of LLMs (increasing model hallucinations) [8], especially for deterministic tasks (e.g., coding), and the diversity of a single model is usually insufficient. Instead of conducting self-consistency, we propose to instruct several LLMs under lower temperatures to generate SQLs and then make votes across the SQLs’ executed results. Such a cross-consistency strategy can not only diversify the SQL queries but also hold LLMs’ performance as low temperatures are set up.
Naive voting across several LLMs
We first use a strong LLM (e.g., GPT4) based on the full schema information to obtain a PreSQL, and parser the PreSQL for schema linking. Then, the full schema prompt is simplified and used to instruct several LLMs, each of which generates a FinSQL with a low temperature. All the FinSQLs and the re-used PreSQL are executed with the SQL database engine and the queried results are obtained. Finally, we check the results of the execution of each SQL and take the majority of the results as the final answer.
Fine-grained voting according to PreSQL complexity
In addition to the naive voting strategy, we further refine the voting rules based on the difficulty of PreSQL. As different LLMs specialize in questions with different complexity grades, putting them together in a voting pool can produce biased results. The complexity of questions is classified into four grades (i.e., easy, medium, hard, and extra) by the abstract syntax trees of PreSQL according to [28], and each grade questions are solved by distinct candidate LLMs for voting. With such fine-grained voting, we can maximize the potential of LLMs and significantly mitigate the voting bias.
4 Experiments
4.1 Experimental Setup
Dataset and Metrics
We evaluate our PET-SQL in Spider [10] benchmark, which is a large-scale cross-domain Text2SQL dataset. It contains 8659 instances in training split and 1034 instances in development split over 200 databases, with non-overlapping databases in each set, and 2147 instances are holdout as the test set across another 34 databases.
The evaluation of Text2SQL performance of methods is conducted by execution accuracy (EX), following the official test-suit [^1] https://github.com/taoyds/test-suite-sql-eval defined in [28]. EX measures the proportion of questions in the evaluation set where the execution results of both the predicted and ground-truth inquiries are identical, relative to the total number of queries.
Evaluated LLMs
Five LLMs are involved to validate the superiority of our framework and we report the best results of our method on the test set, which can surpass rank 2 with about 1% improvement on the Spider leaderboard. The LLMs are: CodeLlama-34B [7], SQLCoder-34B [29], InternLM-70B [30], SenseChat [31], and GPT4-0613 [32]. The first two are specific to coding, especially on SQL, while the others are generic LLMs.
Implemenation details
To decrease the randomness of LLMs output, the temperatures are set extremely low ($10^{-7}$ for all the non-OpenAI LLMs and 0 for GPT4). The max lengths of input and output tokens are 4096 and 200, respectively. In the question skeleton-based demonstration retrieval, we traverse all the database schema and check if the table or column names appear in questions. The matched entities will be masked by <mask> token. Based on the similarities between the sentence embeddings of the target and candidate question skeletons, top-9 similar question-SQL pairs in the demonstration pool are selected and then constructed as the in-context. For cross-consistency, the executed results from the above five LLMs are voting with equal weights.
4.2 Overall Performance
4.2.1 Spider Leaderboard
In Table 1, we report the performance of our method and baselines on the Spider test dataset. Since the leaderboard is closed and rejects new submissions, we tested the performance of our approach offline, with the official test-suit. Thus, it can be concluded that our method achieves the highest execution accuracy among all non-learning-based methods. Specifically, our method surpasses DAIL-SQL by 1% and achieves the best among the open-source methods in the leaderboard  [^2] Since the 1st-rank MiniSeek is not publicly available, we do not engage it in comparisons.. Even with the naive voting strategy, the superiority of our PET-SQL still holds.
Table 1: The overall performance on Spider leaderboard (test set)
Methods EX
RESDSQL-3B + NatSQL [33] 79.9%
C3 + ChatGPT + Zero-Shot [3] 82.3%
DIN-SQL + GPT4 [5] 85.3%
DAIL-SQL + GPT4 + Self-consistency [4] 86.6%
PET-SQL(Naive voting) 86.7%
PET-SQL(Fine-grained voting) 87.6%
4.2.2 Comparison with DAIL-SQL by other foundation LLMs
[…]
Table 2: The comparison between our PET-SQL and DAIL-SQL (9-shot)
Datasets Methods CodeLlama-34B InternLM-70B SQLCoder-34B
Dev DAIL-SQL 0.756 0.742 0.709
PET-SQL(Ours) 0.750 0.750 0.709
Test DAIL-SQL 0.720 0.707 0.697
PET-SQL(Ours) 0.744 0.714 0.741
4.3 The Effect of Prompts
The prompts performance
To evaluate the superiority of the proposed prompt, we compare our few-shot $RE_{p}$ with $CR_{p}$ and $OD_{p}$, which are the recommended prompts in [4]. It should be noted that the selected question-SQL demonstrations in them are the same. They are evaluated on the Spider development and test datasets, separately, and the results are shown in Table 3. The best one is in boldface. It can be seen that our proposed prompt $RE_{p}$ demonstrates superior performance under the zero-shot setting, with roughly 1% to 7% improvements compared to $CR_{p}$. This indicates the versatility of $RE_{p}$, whether for coding-specific or generic LLMs.
Table 3: The performance comparison between different prompt styles (0-shot)
Datasets Prompting Codellama-34B InternLM-70B SQLCoder-34B
Dev $CR_{p}$ 0.685 0.701 0.562
$RE_{p}$ 0.738 0.707 0.657
Test $CR_{p}$ 0.694 0.649 0.631
$OD_{p}$ 0.609 0.673 0.579
$RE_{p}$ 0.717 0.699 0.678
Ablation of $RE_{p}$
Since there are three modifications in $RE_{p}$ compared to $OD_{p}$, that is, optimization rule (OR), cell values references (CV), and foreign key declarations (FK), we perform the ablation experiments on the Spider dev set to provide insights into these components of the proposed objective function specifically. Table 4 shows the ablation results of the $RE_{p}$ on three non-OpenAI LLMs under the zero-shot setting. It is observed that the exclusion of the CV consistently exerts the most significant impact on performance across InternLM and SQLCoder, resulting in about 2.4% and 6.4% decreases in EX. The absence of the OR also demonstrates a notable impact on model performance (about -3%). Interestingly, the FK component exhibits a comparatively smaller influence on $RE_{p}$, even achieving slight improvement compared to the full version. These findings suggest that the CV has the greatest significance, followed by the OR and FK components. This indicates that ensuring the presence of sampled cell values is crucial for LLM generation, and with these references, LLM can better navigate the database and formulate accurate queries. This reduces ambiguity and improves the efficiency of the generated SQL queries.
Table 4: The ablation of the $RE_{P}$ on Spider dev set (0-shot)
CodeLlama-34B InternLM-70B SQLCoder-34B
$RE_{p}$ 73.80% 70.70% 65.70%
        w/o OR 71.50% 69.90% 64.60%
        w/o CV 71.30% 69.10% 61.70%
        w/o FK 66.30% 71.60% 62.30%
4.4 The Effect of Schema Linking
The linked schema is parsed from the PreSQL, which is generated by GPT4-0613 with full schema information. To validate its effect, we introduce the table recall metrics and compare the EX of FinSQLs generated by the LLMs based on simplified and unsimplified schemas, separately.
The table recall metrics are two-fold, and are defined as
  • Totally Equal: $R_{e}={\sum_{i=1}^{N}\mathbf{1}_{e}}/{N}$, where $N$ is the number of instances in the test set. $\mathbf{1}_{e}$ is an indicator function, which returns 1 if the linked tables are exactly the tables that appeared in ground-truth SQL (GT tables) and 0 otherwise. $R_{e}=1$ is the ideal schema linking result, which means all the question-related tables are recalled, and there is no verbose information.
  • Subset: $R_{s}={\sum_{i=1}^{N}\mathbf{1}_{s}}/{N}$. $\mathbf{1}_{s}$ is an indicator function, which returns 1 if GT tables are the subset of the linked tables and 0 otherwise. $R_{s}$ represents the upper limit of the accuracy that LLM can achieve with schema linking. Even if the linked table is not sufficiently simplified, useful tables are still retained in the prompt.
Table 4.4 shows the schema linking (SL for short) recall of GPT4 (i.e., PreSQL). It can be seen that the recall of table linking reaches 94% while the EX of PreSQL is only 85.2%, and the limitations of LLMs may cause the gap. Prompt simplification with linked tables enables LLM to potentially correct the wrong cases in PreSQL. In Table 4.4, we further summarize the average tables mentioned in the prompt and the ground truth (i.e., the average of GT tables). In the full schema (without SL), the average number of tables is 4.89, while in the simplified schema (with SL), it is 1.60 (close to GT). This means that SL significantly simplifies useless table information in prompts.
Table 5: The metrics of SL with GPT4
LLM for SL $R_{e}$ $R_{s}$
GPT4 0.94 0.98
Table 6: The average of tables in prompts
w/o SL w/ SL Ground truth
4.89 1.60 1.57
With the linked tables, we further simplify the prompt and then evaluate the performance of LLMs on the test set with the 9-shot setting. The results are demonstrated in Table 7. The simplified prompt has improved performance across all LLMs, especially for CodeLlama, which validates the effectiveness of the PreSQL-based schema linking.
Table 7: The comparison of the LLMs performance with and without SL (9-shot)
Codellama-34B InternLM-70B SQLCoder-34B GPT4
w/o SL 74.4% 71.4% 74.1% 85.2%
w/ SL 78.2% 75.4% 76.9% 85.5%
4.5 The Effect of Cross-consistency
We evaluate the cross-consistency (CC for short) of the five LLMs mentioned above. In Table 8, we report its performance on dev and test sets. Considering the token cost, we give up the SQL generation by GPT4 on the dev set, and the best result from the remaining LLMs is chosen as the baseline (CodeLlama-34B). For comparison, we also perform self-consistency with CodeLlama-34B (i.e., calling it 5 times) under $temperature=0.5$. On the dev set, compared to the baseline, the gain of self-consistency is trivial, which is only 0.8%, while the cross-consistency with naive voting achieves 7.9% improvements. This proves that cross-consistency is much better than self-consistency. On the test set, it can seen that the fine-grained voting strategy achieves remarkable improvements over the baseline, with 2.1% boosting. Even for the naive version, the improvements are also promising (+1.2%). Compared to naive voting, fine-grained voting can further exploit the potential of LLMs and significantly mitigate the voting bias. For information, self-consistency of five SQLs generated by GPT4 only achieves 0.4% improvements in DAIL-SQL [4]. These findings demonstrate that, despite its power, GPT-4 may generate errors that other LLMs would not. Therefore, introducing several LLMs to complement each other (i.e., cross-consistency) is necessary, which can diversify the results and greatly improve the overall EX.
Table 8: The performance of cross consistency (9-shot)
Methods Dev(w/o GPT4) Test(w/ GPT4)
Baseline (best single LLM) 75.3% 85.5%
Self-consistency 76.1% -
CC(Naive voting) 82.2% 86.7%
CC(Fine-grained voting) - 87.6%
5 Conclusion
This paper presents the PET-SQL framework, which aims to address challenges in Text2SQL tasks by enhancing the prompt and leveraging cross-consistency across LLMs. Our approach achieves 87.6% execution accuracy on the Spider leaderboard. We also propose a PreSQL-based schema linking method to simplify prompt information, improving the efficiency and accuracy of LLMs in generating SQL queries. Overall, the PET-SQL framework demonstrates promising results and opens avenues for further advancements in Text2SQL tasks.